March 2, 2023

From HallCWiki
Jump to: navigation, search

Return to the calendar.

RC Meeting Information

Current RC: Iuliia Skorodumina.

No RC meeting today.


Wednesday Swing (03/01). We started the Swing shift in Controlled Access as Jixie and Michael were still working in the Hall on the Solid detector. Accelerator was recovering from the Maintenance Day. The recovery took longer than expected, so that we saw tune beam only after 8pm. At around 10pm we started receiving production beam and resumed data taking on LD2 target. Swing shift summary: .

Thursday Owl (03/02). We were taking production data overnight at 50 uA. The beam was fairly good. At about 4 am we increased the current ramp rate. The ramp rate was 1 uA/s, which gave 50s ramping for a 50 uA beam, and now it is twice as fast (see This will allow for an increase in effective beam current. According to Dave Meekins, we do not have any target-related restrictions regarding the current ramp rate. Owl shift summary: .

For the current setting (Pm=800 MeV/c), we have collected a total of 323 good counts out of the planned 730.

Current Run Plan

The plan is to continue taking production data for the Pm=800 MeV/c at the highest possible current.